In a genetics class, 6

 students have GREEN eyes, 3

 students have BLUE eyes, and 11

 have HAZEL eyes. If a single student is picked at random, what is the probability that their eyes are GREEN or HAZEL?


1
Expert's answer
2021-10-20T17:26:38-0400

6 students have green eyes = "\frac{6}{20}*100 = 30%"


11 students have hazel eyes "=\frac{11}{20}*100 = 55%"


20 = number of students


So "\frac{6}{20} + \frac{11}{20} = \frac{17}{20}*100 = 85\%"